OpenStudy (anonymous):

Which of the following is not a perfect square trinomial? A. 144 + 12y + y2 B. 100 + 20y + y2 C. 64 – 16y + y2 D. 25 + 10y + y2

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm y^2+20y+100 =(y+10)^2 }\) \(\large\color{midnightblue}{ \rm y^2-16y+64 =(y-8)^2 }\) \(\large\color{midnightblue}{ \rm y^2+10y+25 =(y+5)^2 }\)

OpenStudy (anonymous):

so A?